10. Suppose your university decides to raise tuition fee from Rs. 60,000 per semester to Rs. 80,000 per semester of MBA students. The number of students getting admission is 100. The price elasticity of demand for the students is 0.4. 3 a. Find out the number of students getting admission after the increase in tuition fee. b. Find out the total revenue before and after the increase in tuition fee. c. Explain whether the increase in tuition fee is a wise decision or not on the basis of total revenue findings. 


1
Expert's answer
2022-02-28T11:44:46-0500

a. Find out the number of students getting admission after the increase in tuition fees.

Given:

Ed=0.4,

Q1=100,

P1 = Rs. 60,000,

P2 = Rs. 80,000

The formula to find out elasticity Ed is:

"E_d=\frac{∆Q}{∆P}×\frac{P}{Q};\space where\space ∆Q=Q_2−Q_1 \space\space and\space ∆P=P_2−P_1"


substituting the values given in the formula,

"E_d=−\frac{∆Q}{∆P}×\frac{P_1}{Q_1}"


"0.4=−\frac{∆Q}{(80,000−60,000)}×\frac{60,000}{100}"


"0.4=−\frac{∆Q}{20,000}×600"


"∆Q=−\frac{0.4×20,000}{600} = −13.33"


Since,

"∆Q=Q_2-Q_1\\\Delta Q_2=\Delta Q+Q_1\\Q_2=13.33+100\\Q_2=113.33"


Therefore, the number of students getting admission after an increase in tuition fees will be 113 students.


b) Total revenue (TR) is the product of price and quantity.

"=60,000×100\\=6000000"

Total Revenue before tuition fee rise will be Rs. 6,000,000


After price rise, fees is 80000 and students are 113.33

TR after fee rise

"=80,000×113.33\\=9,066,400"

 

Total Revenue after tuition fee rise will be Rs. 9,066,400.


c. Explain whether the increase in tuition fee is a wise decision or not on the basis of total revenue findings. 

Since after price increase, total revenue is greater than before increase price

(9,066,400 > 6,000,000), increasing tuition fees is a wise decision to increase the revenue of the school.